OpenStudy (anonymous):

A particle moves along the x-axis with a position function x(t). Which position function x(t) is modeled by a velocity function of v(t) = x'(t) = -2 - 2t and an acceleration function of a(t) = v'(t) = -2?

OpenStudy (phi):

x(t) will be the integral of x'(t) in this case x(t) = - 2t - t^2 (+ c ) where c is an unknown constant.

OpenStudy (anonymous):

@phi would it be x(t)=6-2t-t^2

OpenStudy (phi):

that will work. as a test, take the derivative: x' = 0 -2 -2t = -2 - 2t which matches what they told you.

OpenStudy (anonymous):

thanks!

OpenStudy (phi):

yw